Find the solution by substitution:
4x+3y=5
2x-y=-5

Answer by mangopeeler07(462)

4x+3y=5
2x-y=-5

Solve for y in the second equation:
2x-y=-5
-y=-2x-5
y=2x+5

Plug that into the first equation
4x+3y=5
4x+3(2x+5)=5
4x+6x+15=5
10x+15=5
10x=-10
x=-1

Plug that into y=2x+5
y=2(-1)+5
y=-2+5
y=3

So,
x=-1
y=3